# A Survey of Deep Learning-based Radiology Report Generation Using Multimodal Data

### v1

@@ 3.3.1 Image-based feature learning
Previous research primarily utilized CNNs as architectures for extracting image features, however, recently, an increasing number of researchers have opted for the use of Transformers due to their improved performance. Simultaneously, numerous studies proposed novel modules to enhance the model capability. In this section, the model architecture is first introduced, and subsequently enhancement modules are described. These modules include auxiliary tasks, contrastive learning, and memory metrics. The architecture and modules utilized in each paper are outlined in Table LABEL:tab:overall in Appendix A.
CNN and Transformer encoder for feature extraction
The statistics of the architectures used as image feature extractor are shown in Figure 3. In total, forty-six studies extract image features purely based on CNN models. Thirty-four works firstly encode images by CNN and then utilizes the Transformer layers to modify the embeddings. Eight works utilize a pure Transformer architecture to extract image features. Figure 3 shows a clear trend of more studies adopting CNN augmented with Transformer for image feature extraction.
@@
For CNN architecture, two works self-designed CNN models, while the other works built it based on different classical visual models, such as ResNet (He et al., 2016) (42 papers), DenseNet (Huang et al., 2017) (22 papers), VGG (Simonyan and Zisserman, 2014) (5 papers), Faster-RCNN (Ren et al., 2015) (5 papers), Inception-V3 (Szegedy et al., 2016) (1 paper), ResNeXt (Xie et al., 2017) (1 paper), EfficientNet (Tan and Le, 2019) (1 paper), and the Two-Stream Inflated 3D ConvNets (I3D) (Carreira and Zisserman, 2017) (1 paper). Pahwa et al. (2021) modified the HRNet (Sun et al., 2019a), a human pose estimation network. Other three works (Huang et al., 2021a, b, 2022) provided results based on different CNN structures.
@@
To improve model performance, ten works modified the CNN structure by attention modules, which assigned varying degrees of importance (weights) to different parts of the input by learnable parameters, allowing the model to selectively focus on specific regions of an image. Traditional attention mechanisms can be classified into channel-wise (Du et al., 2022; Wang et al., 2022e; Gajbhiye et al., 2022; Pahwa et al., 2021) and spatial-wise (Pahwa et al., 2021; Jia et al., 2021), which allocate different weights to the various channels and spatial positions of the inputs respectively. In addition, Li et al. (2023b) and Wang et al. (2024a) utilized the idea of the class activation map (Zhou et al., 2016) to obtain weights. Yan et al. (2022) initially extracted image patch features, clustered them using an unsupervised method, and then weighted the cluster results. Experimental results show that attention mechanism allows models to pay more attention to the lesions than irrelevant background. With the rise of the Transformers (Vaswani et al., 2017), multi-head attention has become a potent method for information interaction. Wang et al. (2023c) first extracted regions of interest from the frontal view and then employed multi-head attention to fuse information between the frontal and lateral views with the regions. This approach introduced regions of interest into model training to improve model performance.
@@
For Transformer architecture, most of them leverage a standard Transformer encoder, while Li et al. (2023d) argued that aligning images and text posed a challenge due to the continuous nature of images and the discrete nature of text; therefore, they improved the model performance by using a discrete variational autoencoder (Ramesh et al., 2021) to obtain discrete visual tokens. Other works improved the model performance by modifying the self-attention module. Three works (Wang et al., 2022e; Lin et al., 2023; Wang et al., 2023d) added high-order interactions among three inputs of the Transformer attention module. Two works (Miura et al., 2021; Wang et al., 2022e) were inspired by the memory-augmented attention (Cornia et al., 2020) and extended the keys and values with additional plain learnable vectors to record more information. Li et al. (2022b) introduced a learnable parameter in the attention operation. Wang et al. (2023d) modified the encoder by including additional input tokens. These tokens were named ‘expert tokens’ to emulate the “multi-expert joint diagnosis” methodology.
Auxiliary task for feature extraction
@@
Auxiliary tasks aim to provide additional supervision signals to the feature extractor, enabling it to extract information relevant to report generation from images. These tasks mainly include classification (22 papers), graph construction (10 papers), embedding comparison (10 papers), and detection/segmentation (7 papers). Each of them is introduced in detail below.
@@
Classification: The most common auxiliary task used in the reviewed papers is classification referring to assigning images to predefined categories. The predefined categories primarily include medical tags (Gajbhiye et al., 2022; Kaur and Mittal, 2022b; Wang et al., 2022e; Hou et al., 2021b; Du et al., 2022; You et al., 2022; Alfarghaly et al., 2021) and disease labels (Liu et al., 2021e; Zhou et al., 2021; Wang et al., 2022d; Yang et al., 2023; Zhang et al., 2023b; Wang et al., 2024a; Jin et al., 2024; Hou et al., 2021b). The medical tags are from standard medical vocabularies including hundreds of labels, such as anatomical structures and pathological signs. They are provided by manual annotations or auto annotation tools (e.g., NIH MTI web API  [^1] https://ii.nlm.nih.gov/MTI/index.shtml and RadGraph (Jain et al., 2021)). Disease labels are provided by auto annotation tools (e.g., CheXpert (Irvin et al., 2019) and CheXbert (Smit et al., 2020)). Compared to disease labels, medical tags offer a more comprehensive range of information. However, to the best of our knowledge, there is no literature that supports the superiority of medical tags over disease labels. Perhaps due to the extensive scope covered by medical tags, deep learning models face challenges in acquiring such rich knowledge. Zhou et al. (2021) incorporated 32 additional labels for lesion location, size, and shape (e.g., “upper/lower” and “patchy”) into the disease label set, observing a slight improvement in model performance.
@@
Other notable categories used in the reviewed papers include matching status (Li et al., 2023a), local properties (Yang et al., 2021a), report cluster results (Li et al., 2022a), and fix answer categories (Tanwani et al., 2022). Li et al. (2023a) predicted the matching status of a given image-report pair. Yang et al. (2021a) devised localized property labels for breast ultrasound images, such as tumor morphology, to facilitate the identification of properties that are challenging to be discerned in low-resolution images. Li et al. (2022a) first conducted unsupervised clustering on the ground truth report, subsequently utilizing the resultant clusters as labels. This auxiliary task yielded a marked improvement in text generation performance. Tanwani et al. (2022) considered the report generation as a question-answer task, and the classifier was designed for fixed answer categories. In addition, for a detection auxiliary task, classifiers need to be applied to identify attributes of detected regions (e.g., ‘right lung’). This paragraph eschews such cases to circumvent redundancy.
@@
Graph construction: Graph construction aims at introducing prior knowledge into the report generation process. The knowledge graph in this section differs from that in Section 3.1. Here, node features are extracted from images, and edges are defined as parameters in graph convolution networks. In contrast, the node and edge information in the input knowledge graphs is derived from non-image data. A classical method was proposed by Zhang et al. (2020) and yielded promising outcomes. A knowledge graph was constructed firstly based on insights provided by domain experts, where nodes represented major abnormalities and major organs, and bidirectional connections linked nodes that were related to each other. To initialize nodes features, a spatial attention module was introduced after the CNN backbone using 1$\times$1 convolution layers and softmax layers. The number of channels matched the number of nodes. The nodes’ initial embedding was derived as attention-weighted feature maps. Then graph convolution layers were employed to disseminate information throughout the graph, followed by two branches for classification and report generation. First, the classification branch was trained, and subsequently, parameters in both the CNN backbone and the graph convolution layers were frozen, only the report generation decoder was trained. Six works (Liu et al., 2021b, d; Cao et al., 2022; Wang et al., 2022c; Yan, 2022; Zhang et al., 2023b) utilized this method (Zhang et al., 2020). Wang et al. (2022c) expanded the graph (Zhang et al., 2020) by incorporating information from a radiology terms corpus named Radiology Lexicon (RadLex) [^2] http://www.radlex.org/ (Langlotz, 2006). As the number of graph nodes increased, the model performance initially improved, peaked at 40 nodes, and then declined, with a noticeable decrease at 60 nodes. Liu et al. (2021d) constructed a large graph based on the MIMIC-CXR dataset. The nodes represented frequent clinical abnormalities and the edges represented the co-occurrence situation of different abnormalities. In addition, Li et al. (2023b) used the disease prediction results to obtain the node features. The nodes were classification probabilities, with learnable edge weights.
Another graph reconstruction method aims at reconstructing triplets that are in the form of (entity1, relationship, entity2), such as (opacity, suggestive of, infection). Three works (Dalla Serra et al., 2022, 2023b; Li et al., 2022b) firstly predicted the triplets and then generated reports based on them. The experimental results show that using triplets alone for report generation is ineffective; combining them with features extracted from images is necessary for better results.
@@
Embedding comparison: Embedding comparison refers to constraining the consistency of different features in intermediate layers, thereby guiding the learning process. The comparison in reviewed papers is mainly applied between features extracted from images and real reports (Najdenkoska et al., 2021, 2022; Zhou et al., 2021; Yang et al., 2021b; Chen et al., 2022; Wang et al., 2021, 2022d; Yang et al., 2023). Experimental results show that the supervision signals from real text enable extracted visual features carry richer semantic information, facilitating more effective translation into radiology reports. Four works (Wang et al., 2021; Zhou et al., 2021; Wang et al., 2022d; Yang et al., 2023) utilized a triple loss function to compel the image-text paired features to be closer to a latent space than the unpaired ones. Najdenkoska et al. (2021, 2022) inspired by the Auto-Encoding Variational Bayes (Kingma and Welling, 2013). They used real reports to obtain a latent space during training and generated reports based on this space. The image extractors were enabled to capture features from images that closely resemble those found in real reports. Other two works (Yang et al., 2021b; Chen et al., 2022) used the Term Frequency and Inverse Document Frequency (TF-IDF) to extract important information from real reports as supervision signals. TF-IDF is a statistical measure assessing a word’s importance by considering its frequency in a specific document and its rarity across the entire document set.
@@
In addition, to produce reports for abnormalities not seen during training, Sun et al. (2022) initially linearly projected visual features to semantic features, and extracted semantic features of labels by the BioBert model (Lee et al., 2020). Consistent constrain was applied between two similarity: 1) the similarity between pairwise elements in the semantic features from visual features; and 2) the similarity between the semantic features from visual features and the semantic features from labels. Zhang et al. (2023b) integrated semi-supervised learning into report generation using two networks. They first applied different types of noise to an input image to create two variations, which were then fed into the two networks. An auxiliary loss function was employed to ensure consistency in the extracted visual features.
Detection/segmentation: Object detection locates and identifies objects or patterns within an image, focusing on determining their presence and position. Segmentation divides an image into meaningful regions by identifying and separating objects based on specific characteristics. Both processes enhance the model’s understanding of the image by object recognition and region extraction, and can improve the model’s interpretability by linking the detection/segmentation results with generated sentences. The detection/segmentation regions can be anatomical regions (Tanida et al., 2023; Dalla Serra et al., 2023b, a; Wang et al., 2023c; Han et al., 2021; Gu et al., 2024) and abnormal regions (Sun et al., 2022). No literature compares the impact of detection and segmentation tasks on report generation results. However, a publicly available dataset named Chest ImaGenome (see Section 4) offers detection annotations, making them easier to be acquired than segmentation annotations.
In addition, the outputs of auxiliary tasks can provide valuable information such as disease labels, therefore, inputting them into the following generation network is a common choice (Alfarghaly et al., 2021; Hou et al., 2021b; Singh et al., 2021; Yang et al., 2021a; You et al., 2021; Zhou et al., 2021; Du et al., 2022; Jia et al., 2022; Kaur and Mittal, 2022b; Sun et al., 2022; Wang et al., 2022a, e; Yan et al., 2022; You et al., 2022; Tanida et al., 2023; Li et al., 2022b; Jin et al., 2024). For example, Zhou et al. (2021) sent the semantic word embeddings of the predicted findings from the classifier to the report generation decoder.
Contrastive learning for feature extraction
@@
Contrastive learning is a self-supervised learning method to improve the representational capacity of models, which allows models to minimize the distance among positive pairs and maximize it for negative ones. It can be used to train feature extractors (Wang et al., 2022e; Lin et al., 2023; Wu et al., 2022; Wang et al., 2023b). Lin et al. (2023) utilized a classical contrastive learning method named Momentum Contrast (He et al., 2020), where different views or augmented versions of the same image were considered as positive pairs. Another representative contrastive learning work is the Contrastive Language-Image Pre-training (CLIP) model (Radford et al., 2021). It connects textual and visual information by directly training on a vast dataset consisting of image-text pairs. Wang et al. (2022e) directly employed it for image feature extraction and Wu et al. (2022) applied the idea of CLIP to train the feature extractors on the training dataset. Wang et al. (2023b) however argued that previous works treated the entire report as input, overlooking the distinct information contained within individual sentences. This oversight could result in incorrect matching of image-text pairs. Therefore, they proposed phenotype-based contrastive learning. This method involved randomly initializing a set of vectors as phenotypes, allowing sentences and visual embeddings to interact with them, and finally conducting contrastive learning between the processed embeddings. The results outperformed previous contrastive learning methods in report generation.
@@
Contrastive learning also can be part of the training loss (Tanwani et al., 2022; Wang et al., 2022a; Li et al., 2023a; Liu et al., 2023b), and can be applied between visual and textual features (i.e. image-text pairs) (Tanwani et al., 2022; Li et al., 2023a; Liu et al., 2023b), or be applied based on labels, treating samples with shared labels as positives and those without any common labels as negatives (Wang et al., 2022a). Their ablation study demonstrated a significant improvement in results compared to standard contrastive loss (Wang et al., 2022a).
@@
Contrastive attention is another method to utilize contrastive learning. Liu et al. (2021c) designed a contrastive attention model to extract abnormal region features by comparing input samples with normal cases. Similar features shared between the input and normal cases were subtracted from the input image feature, and the remaining feature was then concatenated with the original feature. Song et al. (2022) argued that the contrastive technique (Liu et al., 2021c) did not consider historical information, therefore they proposed a module based on similarity retrieval technique to obtain similar images from the training dataset. The image features were processed by enlarging different features between inputs and the similar retrieved images.
Memory metric for feature extraction
@@
Using a memory metric for image feature extraction assumes the presence of similar features in various medical images. Memory metrics are employed to record and transmit the similarity information during training (Chen et al., 2022; Yan, 2022). Typically, an n×n matrix is randomly initialized, where n represents the number of metric rows. Then, at each training step, the matrix is updated based on the visual features and the previous metrics. The memory metric used in this section is consistent with the metric used in the memory-driven transformer discussed in Section 3.5, with one being applied to images and the other to the generated reports.

### v2

@@ 3.3.1 Image-based feature learning
Previous research primarily utilized CNNs as architectures for extracting image features, however, recently, an increasing number of researchers have opted for the use of Transformers due to their improved performance. Simultaneously, numerous studies proposed novel modules to enhance the model capability. In this section, the model architectures are first introduced and subsequently enhancement modules are described. These modules include auxiliary tasks, contrastive learning, and memory metrics. The architecture and modules utilized in each paper are outlined in Table LABEL:tab:overall in Appendix A.
Architectures for image feature extraction
The statistics of the architectures used as image feature extractors are shown in Figure 5. It shows a clear trend of more studies adopting Transformer-based architectures for image feature extraction. However, unlike the decoder architecture in Section 3.5.1, where the Transformer has entirely replaced the Long Short-Term Memory (LSTM), many studies combine CNNs with Transformers, using CNNs for image processing and feeding the extracted features into the Transformer encoder. However, there is a lack of research comparing this approach with using only CNNs or only Transformers for report generation.
@@
Most studies developed CNN architectures based on classical models, such as the Residual Nets (ResNet) (He et al., 2016) (42 papers) and the Dense Convolutional Network (DenseNet) (Huang et al., 2017) (22 papers). To improve model performance, ten works modified the CNN structure by attention modules, which assigned varying degrees of importance (weights) to different parts of the input by learnable parameters, allowing the model to selectively focus on specific regions of an image, such as small localized lesions. Specifically, the core of CNNs lies in the convolutional kernels, which extracts informative features by processing spatial and channel information within local receptive fields. The introduction of attention mechanisms is based on the assumption that the model requires a mechanism (i.e., attention) to adaptively recalibrate channel or spatial features by explicitly modeling their interdependencies. This enables the model to focus more effectively on the most relevant features. Traditional attention mechanisms can be classified into channel-wise (Du et al., 2022; Wang et al., 2022e; Gajbhiye et al., 2022; Pahwa et al., 2021) and spatial-wise (Pahwa et al., 2021; Jia et al., 2021), which allocate different weights to the various channels and spatial positions of the inputs respectively. In addition, Li et al. (2023c) and Wang et al. (2024a) utilized the idea of the class activation map (Zhou et al., 2016) to obtain weights. Yan et al. (2022) initially extracted image patch features, clustered them using an unsupervised method, and then weighted the cluster results. Experimental results show that the attention mechanism allows models to pay more attention to the lesions than to irrelevant backgrounds. With the rise of the Transformers (Vaswani et al., 2017), multi-head attention has become a potent method for information interaction. In clinical practice, it is widely acknowledged that radiologists report with varying granularity in radiology reports to reflect different degrees of pathological changes (Wang et al., 2023c). Therefore, both global and local information are essential for report generation, with varying importance when generating different content. Dynamic interactions between these elements can facilitate adjustments in their significance, thereby enhancing outcomes. Wang et al. (2023c) extracted regions of interest from the frontal view and then employed multi-head attention to fuse information between the frontal and lateral views with the regions.
@@
For Transformer architectures, most of them leverage a standard Transformer encoder, while two studies (Li et al., 2023e; Lee et al., 2023) noted that aligning images with text is challenging due to the continuous nature of images and the discrete nature of text. To address this, they used pre-trained models (Ramesh et al., 2021; Esser et al., 2021) to generate discrete visual tokens. However, despite the ablation study validation (Li et al., 2023e), the results were still inferior to other report generation methods, likely due to information loss during discretization. Additionally, the pre-trained models were trained on non-medical images, and this domain gap can also impact performance. Other works improved the model performance by modifying the self-attention module. Three works (Wang et al., 2022e; Lin et al., 2023; Wang et al., 2023d) added high-order interactions among the query, key, and value of the Transformer attention module to enhance their interactions for extracting more complex information. Two works (Miura et al., 2021; Wang et al., 2022e) were inspired by the memory-augmented attention (Cornia et al., 2020) and extended the keys and values with additional plain learnable vectors to record more information. Wang et al. (2023d) modified the encoder by including additional input tokens. These tokens were named ‘expert tokens’ to emulate the “multi-expert joint diagnosis” methodology. Ablation studies indicate that these designs enhance the results significantly.
In conclusion, while an increasing number of studies employ Transformer-based architectures for feature extraction, direct comparisons with CNNs remain absent. Most architectural enhancements focus on modifications to attention mechanisms, which have demonstrated effective improvements. Some research explored the discretization of visual features to align with text features; however, it yielded limited success.
Auxiliary loss functions for image feature extraction
@@
Auxiliary loss functions aim to provide additional supervision signals to the feature extractor, enabling it to extract information relevant to report generation from images. Generally, the model optimizes the main and auxiliary loss functions simultaneously through a weighted sum, with the weights reflecting the relative importance of each task. The auxiliary loss functions described in this section are used to constrain the features extracted by the visual encoder. Additionally in this section, we include losses that are contributed to the pre-training phase, such as those from contrastive learning. Although these losses do not work directly with the main loss function, they still serve as auxiliary losses to help build a more effective visual encoder. These tasks mainly include classification (22 papers), graph construction (10 papers), embedding comparison (10 papers), contrastive learning (8 papers), and detection/segmentation (7 papers). Each of them is introduced in detail below.
@@
Classification loss function: The most common auxiliary task used in the reviewed papers is classification referring to assigning images to predefined categories. Ensuring the generated radiology reports contain essential medical facts is crucial. Introducing an auxiliary classification loss helps the model focus on extracting features relevant to these facts during feature extraction. Predefined categories are usually disease labels (Liu et al., 2021e; Zhou et al., 2021; Wang et al., 2022d; Yang et al., 2023; Zhang et al., 2023b; Wang et al., 2024a; Jin et al., 2024; Hou et al., 2021b, 2023), which are provided by auto annotation tools (e.g., CheXpert (Irvin et al., 2019) and CheXbert (Smit et al., 2020a)), typically covering 14 categories of diseases (e.g., atelectasis). For instance, Yang et al. (2023) utilized the CheXpert labeling tool to annotate 14 common chest radiological findings and integrated a classification layer following the visual encoder to predict diseases. The model was optimized by minimizing the binary cross-entropy loss between the true labels and the predicted disease outcomes. Some researchers argue that these categories are insufficient to capture all necessary medical facts for report generation, and thus extend them to include other medical tags (Gajbhiye et al., 2022; Kaur and Mittal, 2022b; Wang et al., 2022e; Hou et al., 2021b; Du et al., 2022; You et al., 2022; Alfarghaly et al., 2021), such as anatomical structures (e.g., abdominal) and pathological signs (e.g., cicatrix). The medical tags are provided by manual annotations, e.g., the Medical Subject Headings annotation (MeSH) for IU-Xray dataset, or auto annotation tools, e.g., the Medical Text Indexer web Application Programming Interface  [^1] https://ii.nlm.nih.gov/MTI/index.shtml and RadGraph (Jain et al., 2021). Compared to disease labels, medical tags offer a more comprehensive range of information. However, to the best of our knowledge, no literature supports the superiority of medical tags over disease labels. Perhaps due to the extensive scope covered by medical tags, deep learning models face challenges in acquiring such rich knowledge. Zhou et al. (2021) incorporated 32 additional labels for lesion location, size, and shape (e.g., “upper/lower” and “patchy”) into the disease label set, observing a slight improvement in model performance.
@@
Other notable categories used in the reviewed papers include matching status (Li et al., 2023b), local properties (Yang et al., 2021a), report cluster results (Li et al., 2022a), and fix answer categories (Tanwani et al., 2022). Li et al. (2023b) proposed a binary classification task to predict the image-report matching status. The image features, extracted by the image encoder, are input into a cross-attention mechanism alongside the reports. The output then passes through a linear layer to predict the matching probability, which is supervised with a cross-entropy loss. The objective of this task is to align the visual features more closely with the text representation. Yang et al. (2021a) devised localized property labels for breast ultrasound images, such as tumor morphology, to facilitate the identification of properties that are challenging to be discerned in low-resolution images. Li et al. (2022a) first conducted unsupervised clustering on the ground truth report, subsequently utilizing the resultant clusters as labels. The study validated the independence of the clustering results to ensure the diversity of the generated themes. The ablation study confirmed the effectiveness of these methods, yet the final results were inferior to those achieved using more direct additional supervisory information, such as disease labels and image-text matching. Tanwani et al. (2022) considered the report generation as a question-answer task, and the classifier was designed for fixed answer categories.
In conclusion, the classification loss function helps the model focus on extracting medically relevant features by providing additional supervision based on medical facts. Other classification tasks include image-text matching, which aligns textual and visual features; local property classification, which aids in identifying details in low-resolution images; and question-answering tasks. Providing more direct disease-related supervision is beneficial for improving results, however, experimental results revealed that overly abundant information may complicate the network’s learning and information extraction, potentially diminishing its effectiveness. In addition, for a detection auxiliary task, classifiers need to be applied to identify attributes of detected regions (e.g., ‘right lung’). This paragraph eschews such cases to circumvent redundancy.
@@
Graph construction loss function: Graph construction aims to integrate prior knowledge into the report generation process, encompassing key organs, diseases, and their interrelations. It enables the network to extract nodes’ features from the image and defines their relationships (i.e., edges) as parameters in graph convolutional networks. A classical method was proposed by Zhang et al. (2020) and yielded promising outcomes. A knowledge graph was constructed firstly based on insights provided by domain experts, where nodes represented major abnormalities and major organs, and bidirectional connections linked nodes that were related to each other. To initialize node features, a spatial attention module was introduced after the CNN backbone using 1$\times$1 convolution layers and softmax layers. The number of channels matched the number of nodes. The nodes’ initial embedding was derived as attention-weighted feature maps. Then graph convolution layers were employed to disseminate information throughout the graph, followed by two branches for classification and report generation. First, the classification branch was trained, and subsequently, parameters in both the CNN backbone and the graph convolution layers were frozen, only the report generation decoder was trained. Six works (Liu et al., 2021b, d; Cao et al., 2022; Wang et al., 2022c; Yan, 2022; Zhang et al., 2023b) utilized this method (Zhang et al., 2020). Wang et al. (2022c) expanded the graph (Zhang et al., 2020) by incorporating information from a radiology terms corpus named Radiology Lexicon (RadLex) [^2] http://www.radlex.org/ (Langlotz, 2006). As the number of graph nodes increased, the model performance initially improved, peaked at 40 nodes, and then declined, with a noticeable decrease at 60 nodes. Liu et al. (2021d) constructed a large graph based on the Medical Information Mart for Intensive Care Chest X-ray (MIMIC-CXR) dataset. The nodes represented frequent clinical abnormalities and the edges represented the co-occurrence situation of different abnormalities. In addition, Li et al. (2023c) used the disease prediction results to obtain the node features. The nodes were classification probabilities, with learnable edge weights.
Another graph reconstruction method aims at reconstructing triplets that are in the form of (entity1, relationship, entity2), such as (opacity, suggestive of, infection). Dalla Serra et al. (2022) proposed a triplet extraction method for chest X-rays, where entities were first extracted using RadGraph (Jain et al., 2021) and then merged via ScispaCy (Neumann et al., 2019) for simplified annotation. Li et al. (2022b) used an AI-accelerated human-in-the-loop method for entity recognition and relation linking in ophthalmic images (Wu et al., 2021b). Three works (Dalla Serra et al., 2022, 2023b; Li et al., 2022b) firstly predicted the triplets and then generated reports based on them. Triplet prediction is supervised by cross-entropy loss in token prediction. The experimental results indicate that integrating the extracted triplet with image features in the subsequent network significantly enhances performance (Li et al., 2022b; Dalla Serra et al., 2023b). The manually constructed graphs mentioned above are highly accurate and pre-defined, allowing the model to mimic human learning from textbooks. However, these graphs are specific to particular diseases; for instance, a graph constructed for chest X-rays cannot be directly applied to retinal diagnostics. Furthermore, the need for manual annotations increases physicians’ workload. To address these limitations, Yang et al. (2023) propose replacing pre-defined graphs with a learnable memory metric (see Section 3.4).
In conclusion, utilizing graph convolutional networks is a mainstream approach for knowledge integration during training, whereas triplet generation is infrequently utilized. Both methods depend on pre-defined information, leading to limited scalability.
@@
Embedding comparison loss function: Embedding comparison refers to constraining the consistency of different features in intermediate layers, thereby guiding the learning process. The comparison in reviewed papers is mainly applied between features extracted from images and real reports (Najdenkoska et al., 2021, 2022; Zhou et al., 2021; Yang et al., 2021b; Chen et al., 2022; Wang et al., 2021, 2022d; Yang et al., 2023), aligning learned visual features with supervised textual features in the same space to facilitate subsequent report generation. Experimental results show that the supervision signals from real text enable the extracted visual features to carry richer semantic information, facilitating more effective translation into radiology reports. Four works (Wang et al., 2021; Zhou et al., 2021; Wang et al., 2022d; Yang et al., 2023) utilized a triple loss function to compel the image-text paired features to be closer to a latent space than the unpaired ones. Najdenkoska et al. (2021, 2022) inspired by the Auto-Encoding Variational Bayes (Kingma and Welling, 2013). They used real reports to obtain a latent space during training and generated reports based on this space. The image extractors were enabled to capture features from images that closely resemble those found in real reports. Other two works (Yang et al., 2021b; Chen et al., 2022) used the Term Frequency and Inverse Document Frequency (TF-IDF) to extract important information from real reports as supervision signals. TF-IDF is a statistical measure assessing a word’s importance by considering its frequency in a specific document and its rarity across the entire document set.
@@
In addition, to produce reports for abnormalities not seen during training, Sun et al. (2022) initially linearly projected visual features to semantic features, and extracted semantic features of labels by the pre-trained Bidirectional Encoder Representations from Transformers for Biomedical Text Mining (BioBert) model (Lee et al., 2020). Consistent constrain was applied between two similarities: 1) the similarity between pairwise elements in the semantic features from visual features; and 2) the similarity between the semantic features from visual features and the semantic features from labels. This operation ensures the similarity between visual features and label features, aligning them while preserving medical facts within the visual features. Zhang et al. (2023b) integrated semi-supervised learning into report generation using teacher-student networks. They first applied different types of noise to an input image to create two variations, which were then fed into the two networks. An auxiliary loss function was employed to ensure consistency in the extracted visual features.
In conclusion, using information from real text reports as supervision signals is highly effective for feature extraction. Visual features that are aligned with the text and rich in semantic content can enhance subsequent report generation. Additionally, embedding-based approaches can inspire future research. For example, if large models demonstrate superior performance, a teacher-student framework could enable smaller models to replicate similar learning outcomes.
@@
Contrastive learning loss function: Contrastive learning is a self-supervised learning method, which allows models to minimize the distance among positive pairs and maximize it for negative ones. Contrastive learning mainly leverages the similarities and differences between samples other than annotations, facilitating the extraction of stronger feature representations. It can be used for image extractor pretraining (Wang et al., 2022e; Lin et al., 2023; Wu et al., 2022; Wang et al., 2023b) or to optimize the model in conjunction with the main loss (i.e., report generation) (Tanwani et al., 2022; Wang et al., 2022a; Li et al., 2023b; Liu et al., 2023c). For pretraining, Lin et al. (2023) utilized a classical contrastive learning method named Momentum Contrast (He et al., 2020), where different views or augmented versions of the same image were considered as positive pairs. Another representative contrastive learning work is the Contrastive Language-Image Pre-training (CLIP) model (Radford et al., 2021). It connects textual and visual information by directly training on a vast dataset consisting of image-text pairs. Wang et al. (2022e) directly employed it for image feature extraction and Wu et al. (2022) applied the idea of CLIP to train the feature extractors on the training dataset. Wang et al. (2023b) however argued that previous works treated th e entire report as input, overlooking the distinct information contained within individual sentences. This oversight could result in incorrect matching of image-text pairs. Therefore, they proposed phenotype-based contrastive learning. This method involved randomly initializing a set of vectors as phenotypes, allowing sentences and visual embeddings to interact with them, and finally conducting contrastive learning between the processed embeddings. The results outperformed previous contrastive learning methods in report generation.
@@
When contrastive learning be part of the training loss, it can be applied between visual and textual features (i.e. image-text pairs) (Tanwani et al., 2022; Li et al., 2023b; Liu et al., 2023c), or be applied based on labels, treating samples with shared labels as positives and those without any common labels as negatives (Wang et al., 2022a). Ablation studies showed a significant improvement when using contrastive learning as an auxiliary loss function.
In conclusion, the additional supervisory signal from contrastive learning arises from variations and similarities among distinct samples. It can be employed for both pre-training and training phases. Given the rich information in medical images and reports, an image pair (or an image-text pair) can form positive pairs for one medical fact while serving as negative pairs for another. Thus, the effective design of positive and negative samples is crucial. Currently, sentence-level matching has demonstrated strong performance for image-text pairs, and utilizing image labels is an effective approach for image pairs.
Detection/segmentation loss function: Object detection locates and identifies objects or patterns within an image, focusing on determining their presence and position. Segmentation divides an image into meaningful regions by identifying and separating objects based on specific characteristics. Both processes improve the model’s comprehension of medical images through object recognition and region extraction. Abnormalities in medical images are often confined to small, localized regions. Supervision by detection or segmentation auxiliary loss enables backbones to yield dense features specific to these regions, rather than the sparse features typically extracted. The detected or segmented regions can be anatomical regions (Tanida et al., 2023; Dalla Serra et al., 2023b, a; Wang et al., 2023c; Han et al., 2021; Gu et al., 2024) and abnormal regions (Sun et al., 2022). Detection is commonly applied to X-ray imaging using the public Chest ImaGenome dataset, while segmentation is only used for spine MRI on a private dataset. The ImaGenome provides atlas-based detection annotations, which offer bounding box labels for anatomical structures (see Section 6). This makes them easier to acquire than segmentation annotations.
Overall, auxiliary loss functions operate in three ways. The primary approach provides disease-related supervisory signals (e.g., classification, graph construction, and detection) to facilitate the model’s identification of specific lesions. The second approach aligns image data with relevant textual information (e.g., embedding comparison), enriching image features with semantic information to make decoding into a report easier. The third approach is contrastive learning, which leverages image similarities and differences to extract more robust features. The careful design of the sample pairs is crucial. In addition, the outputs of auxiliary tasks supervised by auxiliary loss function can provide valuable information such as disease labels, therefore, entering them into the following generation network is a common choice (Alfarghaly et al., 2021; Hou et al., 2021b; Singh et al., 2021; Yang et al., 2021a; You et al., 2021; Zhou et al., 2021; Du et al., 2022; Jia et al., 2022; Kaur and Mittal, 2022b; Sun et al., 2022; Wang et al., 2022a, e; Yan et al., 2022; You et al., 2022; Tanida et al., 2023; Li et al., 2022b; Jin et al., 2024). For example, Zhou et al. (2021) sent the semantic word embeddings of the predicted findings from the classifier to the report generation decoder.
Enhancement modules for image feature extraction
@@
Designing enhancement modules involves creating components to improve the interaction of specific information, thereby enhancing feature extraction efficiency. In the reviewed papers, two main types of modules are employed. Firstly, contrastive attention guides the model’s focus on specific regions. Liu et al. (2021c) designed a contrastive attention model to extract abnormal region features by comparing input samples with normal cases. Similar features shared between the input and normal cases were subtracted from the input image feature, and the remaining feature was then concatenated with the original feature. Song et al. (2022) argued that Liu et al. (2021c) did not consider historical similar cases; therefore, they proposed a module based on similarity retrieval technique to obtain similar images from the training dataset. Unique information in the image is emphasized by enlarging different features between inputs and the retrieved images.
@@
The second is memory metric. Using a memory metric for image feature extraction assumes the presence of similar features in various medical images. Memory metrics are employed to record and transmit the similarity information during training (Chen et al., 2022; Yan, 2022). Typically, an $n\times n$ matrix is randomly initialized, where $n$ represents the number of metric rows. Then, at each training step, the matrix is updated based on the visual features and the previous metrics.